A car started out from Memphis toward Little Rock at the rate of 60 km/h. A second car left from the same point 2 hours later and drove along the same route at 75 km/h. How long did it take the second car to overtake the first car?
Answer by TimothyLamb(4379)

first car distance covered in 2 hours:
s = d/t
d = st
d = 60*2 = 120 km
---
speed of second car relative to the first car:
75 - 60 = 15 km/hr
---
time taken by the second car to overtake the first car:
t = d/s
t = 120/15
t = 8 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php